# theorem]Construction

### v1

@@ 1.2. Percolation thresholds
Whilst weak saturation looks at the smallest possible starting graphs that $H$-percolate, the statistical physics origins of bootstrap percolation motivates the investigation of random initial infection graphs. Balogh, Bollobás, and Morris [12] initiated the study of the threshold probability for when the random graph $G(n,p)$ becomes $H$-percolating. Here $G(n,p)$ denotes the Erdős-Rényi binomial random graph which is obtained by taking every edge on $n$ vertices independently with probability $p=p(n)$. We say an event holds asymptotically almost surely (a.a.s. for short) in $G(n,p)$ if the probability it holds tends to 1 as $n$ tends to infinity. As the property of being $H$-percolating is monotone increasing, the classical result of Bollobás and Thomason [27] gives that there is some threshold $p_{c}=p_{c}(n,H)$, such that if $p=\omega(p_{c})$ then a.a.s. $G(n,p)$ is $H$-percolating whilst if $p=o(p_{c})$ then a.a.s. $G(n,p)$ is not $H$-percolating. Considering Example 1.1, we see that $p_{c}(n,K_{3})$ is precisely the threshold for connectivity in $G(n,p)$, namely $\log n/n$ [41].
@@ 2.1. Weak saturation for cliques
2.2. Percolation thresholds for cliques
@@
where $\lambda(H):=\frac{e_{H}-2}{v_{H}-2}$ and the $o(1)$ in the exponent is hiding explicit poly-logarithmic factors in both the upper bound and lower bounds. For a sanity check we note that $\lambda(K_{3})=1$, as it should be to obtain the connectivity threshold in (2.1). In order to see where the exponent is coming from, consider the simple $K_{k}$-chains from Definition 2.1. Note that if an edge $e\in E(K_{n})$ plays the role of $e_{d}$ for some copy of $F_{d}$ in $G(n,p)$ then certainly $e$ will be infected as one can see that the $K_{k}$-process on $F_{d}$ infects edge $e_{i}$ at time $i$ for $1\leq i\leq d$. Now $n^{-1/\lambda(K_{k})}$ is the threshold for when a fixed edge $e\in E(K_{n})$ plays the role of a non-edge in a copy of $K_{k}^{--}$, the graph obtained from $K_{k}$ by removing a matching of size $2$. Indeed, the expected number of copies of $K_{k}^{--}$ that have $e$ playing the role of a non-edge is $n^{k-2}p^{e(K_{k}^{--})}$ which is constant at $p=n^{-1/\lambda(H)}$. Therefore, if $p$ is significantly above $n^{-1/\lambda(H)}$ (by some poly-logarithmic factors, say), then any $e\in E(K_{n})$ plays the role of a non-edge in $K_{k}^{--}$, and we can string together these copies to start to build some simple $K_{k}$-chain (or rather the resulting graph $F_{d}$) in reverse, that has $e$ as its final edge $e_{d}$. There are in fact many ways to do this and extending these to length $d\geq\log n$ say, we find one in which the final copy of $K_{k}^{--}$ added (which will be the first copy in the chain graph $F_{d}$) is in fact a copy of $K_{k}^{-}$ (that is, there is an extra edge) and so then we have a chain that causes the edge $e$ to be infected. With some more care one can apply the second moment method to show that all edges of $K_{n}$ play the role of $e_{d}$ in some copy of $F_{d}$ in $G(n,p)$ and will hence get infected.
@@ 2.3. Maximum running times for cliques
For larger cliques $K_{k}$, simple chains again give a linear lower bound but there is a key difference with the case of the $K_{4}$-chain. Indeed, for $k\geq 5$ the only edges added in the $K_{5}$-process on $F_{d}$ are the edges $e_{1},\ldots,e_{d}$ with $e_{i}$ added in the $i^{\text{th}}$ time step. This was not the case for $K_{4}$, where the process on $F_{d}$ percolates. The sparse final graph of the process on a simple $K_{k}$-chain gives hope that a long chain of copies of $K_{k}$ can somehow be made to wrap back onto itself whilst maintaining that each edge $e_{i}$ gets infected at time $i$. Bollobás et al. [25] and Matzke [66] both managed to do this and achieve super-linear running times. This birthed the concept of a “chain construction” which has since been the principal idea for deriving lower bounds for running times of $H$-processes. We therefore introduce the relevant definitions in generality allowing us to formalise and extend the ideas above. Moreover, chains as defined below also play a crucial role in deriving upper bounds (see Section 5.3.2).
@@
Setting $G_{0}=G^{\prime}$ to be the starting graph, we show by induction that $G_{i}=G_{0}+\{e_{1},\ldots,e_{i}\}=G_{i-1}+e_{i}$. The base case $i=0$ follows from the definition. For $i\geq 1$, supposing that the claim holds for previous $i$, we have that $H_{i}-e_{i}\subseteq G_{i-1}$ and so certainly $e_{i}$ is added to $G_{i}$. To see that no other edge is added, suppose an edge $f\in G_{i}\setminus G_{i-1}$ completes a copy $F\subseteq G_{i-1}$ of $K_{k}^{-}$ to a copy of $K_{k}$ at time $i$. As $G_{i-1}\subseteq G$ by the induction hypothesis, property 1 gives that $F\subseteq H_{j}$ for some $j\in[\tau]$. As $f\notin G_{i-1}$, we have $j\geq i$. If $j>i$, then $H_{j}\cap G_{i-1}=H_{j}-\{e_{j-1},e_{j}\}$ by the induction hypothesis and condition 1. Therefore there are not enough edges in $H_{j}\cap G_{i-1}$ to contain $F$ and so we must have $j=i$, $F\subseteq H_{i}$ and $f=e_{i}$. ∎
@@
To construct the linearly long simple chains, we start by considering a bipartition of the vertices into two sides which we think of as the left side and a right side. On each side we place a sequence of $2\tau_{0}$ copies of $K_{3}$, with consecutive copies intersecting in a singular vertex. Using these triangles, for any slope $a=0,1,\ldots,\tau_{0}$, we can consider a chain $(H^{a}_{i},e^{a}_{i})_{i\in[\tau_{0}]}$ which matches the triangle in position $j$ on the left to the triangle in position $j+a$ on the right and places a complete bipartite graph between them (see Figure 2). We will take an appropriate collection $A$ of slopes that will define our family of simple chains.
@@
Crucial to the analysis of this chain and establishing the key condition 1 of Lemma 2.3 is the fact that the graph $G:=\cup_{a\in A,i\in[\tau]}H^{a}_{i}$ induces a sequence of triangles on both the left and right side. This in turn forces any copy $F$ of $K_{6}^{-}$ in $G$ to use three vertices on the left and three vertices on the right. Using this, one can show that if $F$ contains edges of the underlying graph $\cup_{i\in[\tau_{0}]}H^{a}_{i}$ for different slopes $a\in A$, then two of these must have difference at most three. We can avoid this by choosing only every fourth integer into $A$.
@@ 3. Proof highlight: $M_{K_{5}}(n)$ lower bound
As with the ladder chains of the previous section, we will first superimpose linearly many simple chains on the same vertex set, each of linear length. Note that each of the chains $(H_{i}^{a},e_{i}^{a})_{i\in[\tau_{0}]}$ there can be obtained from the chain $(H^{0}_{i},e^{0}_{i})_{i\in[\tau_{0}]}$ with slope 0 by leaving the left side fixed and translating the vertices on the right side by $a$ triangles. The main difference for dilation chain constructions will be the way we choose the permutation that “mixes up” the vertices of a simple chain to obtain other simple chains in our family. The vertex set will be labelled by the elements of $\mathbb{Z}_{p}\setminus\{0\}$, which is equipped with nicely interacting additive and multiplicative operations. In the basic simple chain $(H^{1}_{i},e^{1}_{i})_{i\in[\tau_{0}]}$ the copies of $K_{5}$s will be laid out one after the other, following the additive structure of $\mathbb{Z}_{p}\setminus\{0\}$. The family of simple chains $(H^{a}_{i},e^{a}_{i})_{i\in[\tau_{0}]}$ will be created using the multiplicative operation: the vertices of the basic simple chain are multiplied by a constant $a\in\mathbb{Z}_{p}\setminus\{0\}$. We will again find that if we choose the set $A$ of dilation constants $a$ appropriately, then there will be no unwanted copy of $K_{5}^{-}$ when we superimpose the corresponding simple chains. Let us see this idea formally.
@@ 3.1. The construction
Note that in order to have that $(L_{i}^{j},g_{i}^{j})_{i\in[3]}$ is a well-defined simple $K_{5}$-chain, we need that $f^{j}_{\tau_{0}}\in E(L_{1}^{j})$ and $g_{3}^{j}=f^{j+1}_{0}\in E(L_{3}^{j})$ are vertex-disjoint in $W$. We will ensure this when we select the set $A$ of dilations.
@@
In this section, we collect properties of the set $A$ that we will need for our dilation chain construction above to work.
First of all, as we promised above, we make sure that $f^{j}_{\tau_{0}}\cap f^{j+1}_{0}=\varnothing$ for every $j\in[q-1]$, so the $j^{\text{th}}$ linking chain is simple. Furthermore, in order to use Lemma 2.3 for the running time of $(H_{i},e_{i})_{i\in[\tau]}$, we need to establish properties 1 and 1 for it. Both of these depend on certain restrictions about how the edges of different dilation chains interact with each other. At the heart of the proof of all of these lies a general observation, relating the edges of the graph $G^{j}:=\cup_{i\in[\tau_{0}]}H^{j}_{i}$ of the $j^{\text{th}}$ dilation chain to the existence of a congruence modulo $p$.
@@
As $w_{x}w_{y}\in E(G^{j})$, there is some $i\in[\tau_{0}]$ with $w_{x},w_{y}\in V(H^{j}_{i})=\{w_{a_{j}(3i-2)},\ldots,w_{a_{j}(3i+2)}\}$. Therefore, there are some distinct $\ell_{x},\ell_{y}\in\{-2,-1,0,1,2\}$ such that $x=a_{j}(3i+\ell_{x})$ and $y=a_{j}(3i+\ell_{y})$ modulo $p$. Then $x-y=\alpha a_{j}$ modulo $p$, where $\alpha=\ell_{x}-\ell_{y}\neq 0$. ∎
The properties we need for our main proof forbid that edges from two or three different dilation chains form certain small subgraphs in $G$. The above observation translates these subgraphs to certain linear equations modulo $p$, in at most three variables, having a solution from $A$. This leads us to the next theorem, based on the famous construction of Behrend avoiding $3$-term arithmetic progressions. It gives us large subsets $A$ of $\mathbb{Z}_{p}$ which do not have non-trivial solutions to any of the linear equations we will care about.
@@
Note that in particular, by fixing $\alpha_{1}=\alpha_{2}=1$ and $\alpha_{3}=-2$, the set $A$ from Theorem 3.2 avoids three-term arithmetic progressions $x,x+d,x+2d$ with $x,d\in\mathbb{Z}_{p}$ and $d\neq 0$. The original proof of Behrend [19] simply gave sets avoiding these arithmetic progressions but it is well-known [78, 79] that his construction can be adapted to avoid non-trivial solutions to any constant number of linear equations in at most three variables.
Distinct dilation chains are edge-disjoint. To guarantee 1 we need to make sure that $G^{j}$ and $G^{j^{\prime}}$ do not share an edge if $j\neq j^{\prime}$. If there was some edge $w_{x}w_{y}\in G_{j}\cap G_{j^{\prime}}$ then by Observation 3.1 we have that there are some $\alpha,\alpha^{\prime}\in\mathbb{Z}\setminus\{0\}$ with $|\alpha|,|\alpha^{\prime}|\leq 4$ such that $\alpha a_{j}-\alpha^{\prime}a_{j^{\prime}}=0$ modulo $p$. This contradicts our choice of $A$ from Theorem 3.2, since $a_{j},a_{j^{\prime}}\in A$ and $a_{j}\neq a_{j^{\prime}}$.
Corollary 3.3.
@@
Preventing “unwanted” copies of $K_{5}^{-}$. Condition 1 forbids that a copy $F$ of $K_{5}^{-}$ in $G$ shares an edge with more than one $H_{i}$. The most problematic part of showing this turns out to be if $F$ shares an edge with more than one dilation chain. Here is where we use the full power of the properties of our set $A$ from Theorem 3.2.
@@
modulo $p$. Summing these equations gives that $\sum_{i=1}^{3}\alpha_{i}a_{i}=0$ modulo $p$ and thus, as all the $\alpha_{i}$ are non-zero and $A$ satisfies the conclusion of Theorem 3.2, we have that $a_{j_{1}}=a_{j_{2}}=a_{j_{3}}$, contradicting that $F^{\prime}$ is non-monochromatic. ∎
Linking chains are simple. Finally, we also rectify what we left open at the end of the definition of linking chains: to ensure that they are simple. Namely, we need to see that for every $j\in[q-1]$ the edge $f^{j}_{\tau_{0}}$ and the edge $f^{j+1}_{0}$ are disjoint. The vertices of $f^{j+1}_{0}$ are $w_{a_{j+1}}$ and $w_{2a_{j+1}}$ and the vertices of $f^{j}_{\tau_{0}}$ are among $w_{a_{j}(p-\alpha)}$, where $\alpha\in[4]$. If two of these were to coincide then there was $\alpha^{\prime}\in[2]$ and $\alpha\in[4]$ such that $\alpha^{\prime}a_{j+1}-\alpha a_{j}=0$ modulo $p$, which contradicts the property of our set $A$.
@@
3.3. Finishing the proof
Finally we show that for appropriate choice of $p$ and $A$, the $K_{5}$-chain $(H_{i},e_{i})_{i\in[\tau]}$ satisfies the conditions of Lemma 2.3, hence the starting graph $G^{\prime}:=(\cup_{i\in[\tau]}H_{i})-\{e_{1},\ldots,e_{\tau}\}$ has running time $\tau=n^{2-o(1)}$. To this end, for $n\in\mathbb{N}$ sufficiently large, we let $p\in\mathbb{N}$ be a prime such that $n/4\leq p\leq n/2$, which exists by Bertrand’s postulate/Chebyshev’s theorem, and fix $\tau_{0}:=\lfloor\frac{p-3}{3}\rfloor$ as before. Let $A:=\{a_{1},\ldots,a_{q}\}\subseteq\mathbb{Z}_{p}\setminus\{0\}$ with $q\geq p^{1-C^{\prime}/\sqrt{\log p}}$ for some $C^{\prime}>0$ be the set from Theorem 3.2. Then for an appropriate $C>0$ we have

### v2

@@ 1.2. Percolation thresholds
Whilst weak saturation looks at the smallest possible starting graphs that $H$-percolate, the statistical physics origins of bootstrap percolation motivates the investigation of random initial infection graphs. Balogh, Bollobás, and Morris [12] initiated the study of the threshold probability for when the random graph $G(n,p)$ becomes $H$-percolating. Here $G(n,p)$ denotes the Erdős-Rényi binomial random graph which is obtained by taking every edge on $n$ vertices independently with probability $p=p(n)$. We say an event holds asymptotically almost surely ($a.$s. for short) in $G(n,p)$ if the probability it holds tends to 1 as $n$ tends to infinity. As the property of being $H$-percolating is monotone increasing, the classical result of Bollobás and Thomason [27] gives that there is some threshold $p_{c}=p_{c}(n,H)$, such that if $p=\omega(p_{c})$ then $a.$s. $G(n,p)$ is $H$-percolating whilst if $p=o(p_{c})$ then $a.$s. $G(n,p)$ is not $H$-percolating. Considering Example 1.1, we see that $p_{c}(n,K_{3})$ is precisely the threshold for connectivity in $G(n,p)$, namely $\log n/n$ [41].
@@ 2.1. Weak saturation for cliques
$2.$ Percolation thresholds for cliques
@@
where $\lambda(H):=\frac{e_{H}-2}{v_{H}-2}$ and the $o(1)$ in the exponent is hiding explicit poly-logarithmic factors in both the upper bound and lower bounds. For a sanity check we note that $\lambda(K_{3})=1$, as it should be to obtain the connectivity threshold in (2.1). In order to see where the exponent is coming from, consider the simple $K_{k}$-chains from Definition 2.1. Note that if an edge $e\in E(K_{n})$ plays the role of $e_{d}$ for some copy of $F_{d}$ in $G(n,p)$ then certainly $e$ will be infected as one can see that the $K_{k}$-process on $F_{d}$ infects edge $e_{i}$ at time $i$ for $1\leq i\leq d$. Now $n^{-1/\lambda(K_{k})}$ is the threshold for when a fixed edge $e\in E(K_{n})$ plays the role of a non-edge in a copy of $K_{k}^{--}$, the graph obtained from $K_{k}$ by removing a matching of size 22. Indeed, the expected number of copies of $K_{k}^{--}$ that have $e$ playing the role of a non-edge is $n^{k-2}p^{e(K_{k}^{--})}$ which is constant at $p=n^{-1/\lambda(H)}$. Therefore, if $p$ is significantly above $n^{-1/\lambda(H)}$ (by some poly-logarithmic factors, say), then any $e\in E(K_{n})$ plays the role of a non-edge in $K_{k}^{--}$, and we can string together these copies to start to build some simple $K_{k}$-chain (or rather the resulting graph $F_{d}$) in reverse, that has $e$ as its final edge $e_{d}$. There are in fact many ways to do this and extending these to length $d\geq\log n$ say, we find one in which the final copy of $K_{k}^{--}$ added (which will be the first copy in the chain graph $F_{d}$) is in fact a copy of $K_{k}^{-}$ (that is, there is an extra edge) and so then we have a chain that causes the edge $e$ to be infected. With some more care one can apply the second moment method to show that all edges of $K_{n}$ play the role of $e_{d}$ in some copy of $F_{d}$ in $G(n,p)$ and will hence get infected.
@@ 2.3. Maximum running times for cliques
For larger cliques $K_{k}$, simple chains again give a linear lower bound but there is a key difference with the case of the $K_{4}$-chain. Indeed, for $k\geq 5$ the only edges added in the $K_{5}$-process on $F_{d}$ are the edges $e_{1},\ldots,e_{d}$ with $e_{i}$ added in the $i^{\text{th}}$ time step. This was not the case for $K_{4}$, where the process on $F_{d}$ percolates. The sparse final graph of the process on a simple $K_{k}$-chain gives hope that a long chain of copies of $K_{k}$ can somehow be made to wrap back onto itself whilst maintaining that each edge $e_{i}$ gets infected at time ii. Bollobás et al. [25] and Matzke [66] both managed to do this and achieve super-linear running times. This birthed the concept of a “chain construction” which has since been the principal idea for deriving lower bounds for running times of $H$-processes. We therefore introduce the relevant definitions in generality allowing us to formalise and extend the ideas above. Moreover, chains as defined below also play a crucial role in deriving upper bounds (see Section 5.3.2).
@@
Setting $G_{0}=G^{\prime}$ to be the starting graph, we show by induction that $G_{i}=G_{0}+\{e_{1},\ldots,e_{i}\}=G_{i-1}+e_{i}$. The base case $i=0$ follows from the definition. For $i\geq 1$, supposing that the claim holds for previous $i$, we have that $H_{i}-e_{i}\subseteq G_{i-1}$ and so certainly $e_{i}$ is added to $G_{i}$. To see that no other edge is added, suppose an edge $f\in G_{i}\setminus G_{i-1}$ completes a copy $F\subseteq G_{i-1}$ of $K_{k}^{-}$ to a copy of $K_{k}$ at time ii. As $G_{i-1}\subseteq G$ by the induction hypothesis, property 1 gives that $F\subseteq H_{j}$ for some $j\in[\tau]$. As $f\notin G_{i-1}$, we have $j\geq i$. If $j>i$, then $H_{j}\cap G_{i-1}=H_{j}-\{e_{j-1},e_{j}\}$ by the induction hypothesis and condition 1. Therefore there are not enough edges in $H_{j}\cap G_{i-1}$ to contain $F$ and so we must have $j=i$, $F\subseteq H_{i}$ and $f=e_{i}$. ∎
@@
To construct the linearly long simple chains, we start by considering a bipartition of the vertices into two sides which we think of as the left side and a right side. On each side we place a sequence of $2\tau_{0}$ copies of $K_{3}$, with consecutive copies intersecting in a singular vertex. Using these triangles, for any slope $a=0,1,\ldots,\tau_{0}$, we can consider a chain $(H^{a}_{i},e^{a}_{i})_{i\in[\tau_{0}]}$ which matches the triangle in position $j$ on the left to the triangle in position $j+a$ on the right and places a complete bipartite graph between them (see Figure 2). We will take an appropriate collection AA of slopes that will define our family of simple chains.
@@
Crucial to the analysis of this chain and establishing the key condition 1 of Lemma 2.3 is the fact that the graph $G:=\cup_{a\in A,i\in[\tau]}H^{a}_{i}$ induces a sequence of triangles on both the left and right side. This in turn forces any copy $F$ of $K_{6}^{-}$ in $G$ to use three vertices on the left and three vertices on the right. Using this, one can show that if $F$ contains edges of the underlying graph $\cup_{i\in[\tau_{0}]}H^{a}_{i}$ for different slopes $a\in A$, then two of these must have difference at most three. We can avoid this by choosing only every fourth integer into AA.
@@ 3. Proof highlight: $M_{K_{5}}(n)$ lower bound
As with the ladder chains of the previous section, we will first superimpose linearly many simple chains on the same vertex set, each of linear length. Note that each of the chains $(H_{i}^{a},e_{i}^{a})_{i\in[\tau_{0}]}$ there can be obtained from the chain $(H^{0}_{i},e^{0}_{i})_{i\in[\tau_{0}]}$ with slope 0 by leaving the left side fixed and translating the vertices on the right side by $a$ triangles. The main difference for dilation chain constructions will be the way we choose the permutation that “mixes up” the vertices of a simple chain to obtain other simple chains in our family. The vertex set will be labelled by the elements of $\mathbb{Z}_{p}\setminus\{0\}$, which is equipped with nicely interacting additive and multiplicative operations. In the basic simple chain $(H^{1}_{i},e^{1}_{i})_{i\in[\tau_{0}]}$ the copies of $K_{5}$s will be laid out one after the other, following the additive structure of $\mathbb{Z}_{p}\setminus\{0\}$. The family of simple chains $(H^{a}_{i},e^{a}_{i})_{i\in[\tau_{0}]}$ will be created using the multiplicative operation: the vertices of the basic simple chain are multiplied by a constant $a\in\mathbb{Z}_{p}\setminus\{0\}$. We will again find that if we choose the set AA of dilation constants $a$ appropriately, then there will be no unwanted copy of $K_{5}^{-}$ when we superimpose the corresponding simple chains. Let us see this idea formally.
@@ 3.1. The construction
Note that in order to have that $(L_{i}^{j},g_{i}^{j})_{i\in[3]}$ is a well-defined simple $K_{5}$-chain, we need that $f^{j}_{\tau_{0}}\in E(L_{1}^{j})$ and $g_{3}^{j}=f^{j+1}_{0}\in E(L_{3}^{j})$ are vertex-disjoint in $W$. We will ensure this when we select the set AA of dilations.
@@
In this section, we collect properties of the set AA that we will need for our dilation chain construction above to work.
First of all, as we promised above, we make sure that $f^{j}_{\tau_{0}}\cap f^{j+1}_{0}=\varnothing$ for every $j\in[q-1]$, so the $j^{\text{th}}$ linking chain is simple. Furthermore, in order to use Lemma 2.3 for the running time of $(H_{i},e_{i})_{i\in[\tau]}$, we need to establish properties 1 and 1 for it. Both of these depend on certain restrictions about how the edges of different dilation chains interact with each other. At the heart of the proof of all of these lies a general observation, relating the edges of the graph $G^{j}:=\cup_{i\in[\tau_{0}]}H^{j}_{i}$ of the $j^{\text{th}}$ dilation chain to the existence of a congruence modulo pp.
@@
As $w_{x}w_{y}\in E(G^{j})$, there is some $i\in[\tau_{0}]$ with $w_{x},w_{y}\in V(H^{j}_{i})=\{w_{a_{j}(3i-2)},\ldots,w_{a_{j}(3i+2)}\}$. Therefore, there are some distinct $\ell_{x},\ell_{y}\in\{-2,-1,0,1,2\}$ such that $x=a_{j}(3i+\ell_{x})$ and $y=a_{j}(3i+\ell_{y})$ modulo pp. Then $x-y=\alpha a_{j}$ modulo $p$, where $\alpha=\ell_{x}-\ell_{y}\neq 0$. ∎
The properties we need for our main proof forbid that edges from two or three different dilation chains form certain small subgraphs in $G$. The above observation translates these subgraphs to certain linear equations modulo $p$, in at most three variables, having a solution from AA. This leads us to the next theorem, based on the famous construction of Behrend avoiding 33-term arithmetic progressions. It gives us large subsets AA of $\mathbb{Z}_{p}$ which do not have non-trivial solutions to any of the linear equations we will care about.
@@
Note that in particular, by fixing $\alpha_{1}=\alpha_{2}=1$ and $\alpha_{3}=-2$, the set AA from Theorem 3.2 avoids three-term arithmetic progressions $x,x+d,x+2d$ with $x,d\in\mathbb{Z}_{p}$ and $d\neq 0$. The original proof of Behrend [19] simply gave sets avoiding these arithmetic progressions but it is well-known [78, 79] that his construction can be adapted to avoid non-trivial solutions to any constant number of linear equations in at most three variables.
Distinct dilation chains are edge-disjoint. To guarantee 1 we need to make sure that $G^{j}$ and $G^{j^{\prime}}$ do not share an edge if $j\neq j^{\prime}$. If there was some edge $w_{x}w_{y}\in G_{j}\cap G_{j^{\prime}}$ then by Observation 3.1 we have that there are some $\alpha,\alpha^{\prime}\in\mathbb{Z}\setminus\{0\}$ with $|\alpha|,|\alpha^{\prime}|\leq 4$ such that $\alpha a_{j}-\alpha^{\prime}a_{j^{\prime}}=0$ modulo pp. This contradicts our choice of AA from Theorem 3.2, since $a_{j},a_{j^{\prime}}\in A$ and $a_{j}\neq a_{j^{\prime}}$.
Corollary $3.$
@@
Preventing “unwanted” copies of $K_{5}^{-}$. Condition 1 forbids that a copy $F$ of $K_{5}^{-}$ in $G$ shares an edge with more than one $H_{i}$. The most problematic part of showing this turns out to be if $F$ shares an edge with more than one dilation chain. Here is where we use the full power of the properties of our set AA from Theorem 3.2.
@@
modulo pp. Summing these equations gives that $\sum_{i=1}^{3}\alpha_{i}a_{i}=0$ modulo $p$ and thus, as all the $\alpha_{i}$ are non-zero and AA satisfies the conclusion of Theorem 3.2, we have that $a_{j_{1}}=a_{j_{2}}=a_{j_{3}}$, contradicting that $F^{\prime}$ is non-monochromatic. ∎
Linking chains are simple. Finally, we also rectify what we left open at the end of the definition of linking chains: to ensure that they are simple. Namely, we need to see that for every $j\in[q-1]$ the edge $f^{j}_{\tau_{0}}$ and the edge $f^{j+1}_{0}$ are disjoint. The vertices of $f^{j+1}_{0}$ are $w_{a_{j+1}}$ and $w_{2a_{j+1}}$ and the vertices of $f^{j}_{\tau_{0}}$ are among $w_{a_{j}(p-\alpha)}$, where $\alpha\in[4]$. If two of these were to coincide then there was $\alpha^{\prime}\in[2]$ and $\alpha\in[4]$ such that $\alpha^{\prime}a_{j+1}-\alpha a_{j}=0$ modulo $p$, which contradicts the property of our set AA.
@@
$3.$ Finishing the proof
Finally we show that for appropriate choice of $p$ and AA, the $K_{5}$-chain $(H_{i},e_{i})_{i\in[\tau]}$ satisfies the conditions of Lemma 2.3, hence the starting graph $G^{\prime}:=(\cup_{i\in[\tau]}H_{i})-\{e_{1},\ldots,e_{\tau}\}$ has running time $\tau=n^{2-o(1)}$. To this end, for $n\in\mathbb{N}$ sufficiently large, we let $p\in\mathbb{N}$ be a prime such that $n/4\leq p\leq n/2$, which exists by Bertrand’s postulate/Chebyshev’s theorem, and fix $\tau_{0}:=\lfloor\frac{p-3}{3}\rfloor$ as before. Let $A:=\{a_{1},\ldots,a_{q}\}\subseteq\mathbb{Z}_{p}\setminus\{0\}$ with $q\geq p^{1-C^{\prime}/\sqrt{\log p}}$ for some $C^{\prime}>0$ be the set from Theorem 3.2. Then for an appropriate $C>0$ we have